OpenStudy (anonymous):

Simplify. square root of 7 over fourth root of 7 7 to the power of negative 3 over 4 7 to the power of 1 over 4 7 to the power of 3 over 4 7

OpenStudy (johnweldon1993):

\[\large \frac{\sqrt{7}}{\sqrt[4]{7}} = \frac{7^{1/2}}{7^{1/4}} = 7^{1/2 - 1/4} = 7^{1/4} = \sqrt[4]{7} \]

OpenStudy (anonymous):

@johnweldon1993 Can you help me with another one?

OpenStudy (johnweldon1993):

sure why not

OpenStudy (anonymous):

The function below shows the relationship between the length of each side of a square (y) and the area of the square (x + 3): y equals square root of the whole of x plus 3 Which graph best shows the function? Graph of y equals square root of the whole of x plus 3 graph of y equals cube root of the whole of x plus 3 graph of y equals square root of the whole of x minus 3 graph of y equals cube root of the whole of x minus 3

OpenStudy (anonymous):

@johnweldon1993